How many total atoms are in 0.270 g of P2O5?
1
Expert's answer
2015-10-28T11:58:15-0400
Each P2O5 molecule consists of 7 atoms (2P and 5O). Let's first calculate the amount of P2O5 molecules.
N(P2O5) = NA*n(P2O5), where n(P2O5) = m(P2O5)/M(P2O5) and NA is Avogadro constant and 6.022x1023 molecules/mol
N(P2O5) = NA*m(P2O5)/M(P2O5) = 6.022x1023*0.27/142 = 1.145x1021 molecules
So the total number of atoms is:
N(Atoms) = 7* N(P2O5) = 7*1.145x1021 = 8.015x1021 molecules

Answer : Total number of atoms is 8.015x1021 molecules.